OpenStudy (anonymous):

put x^2+y^2-2x+4y+1=0 into the standard form for circles. Show work.

OpenStudy (anonymous):

(x-1)^2 + (y + 2)^2 = 2^2

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

add and subtract 4 and 1

OpenStudy (lalaly):

\[x^2-2x+y^2-4y-1=0\]\[(x-1)^2-1+(y+2)^2-4+1=0\] \[(x-1)^2+(y+2)^2=4\]

OpenStudy (anonymous):

Like lana

OpenStudy (mimi_x3):

(x-1)^2 - 1^2 + (y+2)^2 - 2^2 + 1 = 0 (x-1)^2 + (y+2)^2 -4 = 0 C(1,-2) R = 2

OpenStudy (anonymous):

i dont get how you guys got these answers

OpenStudy (mimi_x3):

Hm, using the formula (x-h)^2 + (y+k)^2 = r^2

OpenStudy (anonymous):

how do you find h and k tho?

OpenStudy (lalaly):

i completed the square first for x^-2x then for y^2+4y

OpenStudy (anonymous):

what did you get when you compleated the squares? and how?

OpenStudy (lalaly):

\[x^2-2x = (x-\frac{2}{2})^2-(\frac{2}{2})^2\] \[y^2+4y= (y+\frac{4}{2})^2-(\frac{4}{2})^2\]

OpenStudy (anonymous):

do you subtract on the opposite side, if not, then whats the point of subtracting?

OpenStudy (mimi_x3):

Use the formula x^2 + bx = (x+b/2)^2 - (b/2)^2 x^2 - bx = (x-b/2)^2 - (b/2)^2

OpenStudy (lalaly):

this is how i complete the square \[x^2+bx= (x+\frac{b}{2})^2-(\frac{b}{2})^2\]

OpenStudy (anonymous):

so rather than subtracting (b/2)^2, can i just add it to the other side?

OpenStudy (lalaly):

yeah u can

OpenStudy (anonymous):

so suppose we're using this question (which im still confused on), it becomes (x+1)^2+(y+2)^2=4

OpenStudy (anonymous):

after that do i expand the parenthetical's?

OpenStudy (mimi_x3):

Expand!? (x+1)^2+(y+2)^2=4 thats already the standard form of the circle

OpenStudy (anonymous):

that become the standard form then? so center would be (-1,-2)?

OpenStudy (mimi_x3):

Yeah for that circle the centre is (-1,-2) and the radius is 2

OpenStudy (anonymous):

ohh, thanks a whole bunch! i get it now!... i think...

OpenStudy (mimi_x3):

lols "i think" tell me if you don't get it i can explain further

OpenStudy (anonymous):

im gonna try another similar question on my hw and then check back in here

OpenStudy (anonymous):

so, can you tell me how to solve for

OpenStudy (lalaly):

center is (1,-2) not (-1,-2)

OpenStudy (anonymous):

(x+1)^2+ [(2x+1)+2]^2=4?

OpenStudy (mimi_x3):

Um , for (x+1)^2+(y+2)^2=4 the C(-1,-2) Xd

OpenStudy (mimi_x3):

xD*

OpenStudy (anonymous):

wait, now how did you get (1,-2) instead of (-1,-2)?

OpenStudy (mimi_x3):

No, she was looking at your previous question xD

OpenStudy (lalaly):

oh lol im sleepy

OpenStudy (anonymous):

the one up top?

OpenStudy (lalaly):

yeh sory ,, didn know its a new question nw

OpenStudy (mimi_x3):

(x+1)^2+ [(2x+1)+2]^2=4? Is that right ?

OpenStudy (anonymous):

wow, no, i actually need the answer to the original question

OpenStudy (lalaly):

ohlol then am not sleepy

OpenStudy (mimi_x3):

I already gave you the answer to the orginal one xD

OpenStudy (anonymous):

but yes, what im trying to do in my hw problem is disprove line y=2x+1 intersecting with circle X^2+y^2-2x+4y+1=0

OpenStudy (anonymous):

so im a little confused again, did i set the second part up right?

OpenStudy (mimi_x3):

disprove!? what do you mean by that

OpenStudy (anonymous):

show that (said line) does not intersect (said equation)

OpenStudy (anonymous):

so did i set it up right?

OpenStudy (mimi_x3):

Set up?

OpenStudy (anonymous):

so im trying to disprove line y=2x+1 intersecting with circle X^2+y^2-2x+4y+1=0 we got down to the standard form of (x+1)^2+(y+2)^2=4 and y=2x+1, when i plug it in is it supposed to look like this: (x+1)^2+[(2x+1)+2]^2=4?

OpenStudy (lalaly):

yeh thats right

OpenStudy (anonymous):

ok, i think i got the rest then... hopfully.

OpenStudy (anonymous):

wait,

OpenStudy (lalaly):

anyways if u need any help am here

OpenStudy (anonymous):

what i ment to type was, how am i supposed to expand [(2x+1)+2]^2?

OpenStudy (anonymous):

sry, didnt post

OpenStudy (lalaly):

\[(2x+3)(2x+3)= 4x^2+12x+9\]

OpenStudy (anonymous):

ok, so i add the 2 before the i square?

OpenStudy (lalaly):

yeh coz its added

OpenStudy (anonymous):

ok, thanks alot

OpenStudy (lalaly):

ur welcome=)